hello i need help on this question.
a car driven 240 miles in 12 hours. what is the rate in miles per hour? in hours per mile? i got the miles per hour is 20. i just cant figure out how to find hour per mile.
jim

Found 2 solutions by rfer, checkley79:
Answer by rfer(16322)

240/12=20mph
1/20=0.05 hrs per mi

Answer by checkley79(3341)

D=RT
R=D/T
1/R=T/D
1/20=12/240
1/20=1/20 OR .05 HOURS (OR .05*60=3 MINUTES PER MILE).